There are birds on two trees.If a bird flew from first to second tree then there are equal number of birds on each tree.However if a bird flew from second to first tree then the birds on first tree are double than the birds on second tree.How many birds are on each tree?
Answer by mananth(16949)

let there be x birds on tree 1
y birds on tree2
one flies from tree 1
x-1= y+1
x-y=2............1
One flies from tree2
x+1 = 2(y-1)
x+1 =2y-2
x-2y=-3...............2
multiply equation (1) by -1
-x+y =-2
add to (2)
x-2y=-3


-y = -5
y=5
plug y =5 in any equation
we get x=7
tree 1 has 7 birds
& tree 2 has 5 birds